# Topological Big Bangs: Reflection, Itty-Bitty Blenders, and Eternal Trumpets

### v2

@@ 4 Beyond Reflection
We have so far taken some care to emphasize the qualifier reflective in our discussion of topological big bangs. This is because one might imagine that it is possible to replace the singularity of cosmological models with some topology in a manner qualitatively different than Definition 2.1, but still consistent with the broader idea of a topological big bang. We exhibit that this is indeed the case with an explicit example.
@@ 4.1 The Itty-Bitty Blender
Note that the coordinate vector field $\partial_{\beta}$ is always spacelike, so the last $S^{1}$ factor does not interestingly affect the overall structure, and we suppress this factor in visualizations and discussion. In contrast, $\partial_{r}$ and $\partial_{\alpha}$ exchange causal character at $r=1$. For $r<1$, $\partial_{r}$ is spacelike and $\partial_{\alpha}$ is timelike: in this range, one should envision that $(r,\theta,\alpha)$ parameterizes a solid toroidal region with $r$ the radial coordinate in the circular cross section at azimuthal angle $\alpha$, and with $\theta$ the angular coordinate around this cross section (see Figure 4).
@@
As in the case of reflective topological big bangs, the causal pathologies of this spacetime (closed timelike curves) are localized to the very early universe. One may still make sense of a classical initial value problem for the future of any spatial $T^{3}$ slice outside the blender (at any $r_{0}>1$), and an observer in the present has no way of accessing a blender cosmology’s closed timelike curves. It is interesting to note that there is no need for inflation to resolve the horizon problem in this picture: the eternal mixing accommodated by the blender ensures that the whole of spacetime is thoroughly causally embroiled, as every inextendible causal curve emanated from the blender.
Of course, we do not attempt here to replicate other successes of inflationary perturbative cosmology (see [1]) or claim that the present model is a genuinely compelling alternative without much more work. Indeed, the metric (4.1) is somewhat contrived and was composed with no thought toward physical energy conditions– in fact, a computation shows that it satisfies none of the standard energy conditions in the $r<1$ region. It is an interesting question whether a qualitatively similar picture can be attained under, say, the dominant energy condition or via a physically motivated matter field; we discuss interplay with the singularity theorems below. Instead, this example, as the whole of this work, is a proof of concept that there exists a rich vein of possibilities for the nature of the early universe beyond the standard discourse.
@@ 4.2 The Blender Unwound: An Eternal Trumpet
Reflective topological big bangs and the Itty-Bitty Blender have an additional out through their causal pathologies– they are clearly not globally hyperbolic, so the cosmological spatial slices $\Sigma_{t}$ are not Cauchy surfaces. While it is well known, of course, that sufficiently egregious causal pathologies allow one to work around most singularity theorems, these examples are interesting in that they illustrate that such pathologies can be entirely localized to the universe’s “earliest” moments. It may appear conceivable, then, that one could write down a metric respecting the strong energy condition while having the same qualitative structure as the Itty-Bitty Blender. This is not the case for the blender as we have presented it, with cosmological spatial slices being copies of $T^{3}$ (and hence compact), due to the following variant of Hawking’s theorem (Theorem 55B in [24]; a relevant result is also established in [21]):
@@
Any blender-type construction on $\mathbb{R}^{2}\times T^{2}$, then, cannot respect the strong energy condition without introducing timelike incompleteness. One could reasonably accept this and seek a metric only satisfying an independent (and more physical) energy condition, such as the dominant energy condition– there appears to be no obstruction to this in the literature:
Question 4.3.
@@
It would be very interesting for considerations of what constitutes a viable model of cosmology if the answer to this question were affirmative. Another approach to working around the constraints of Theorems 4.1 and 4.2 is to “unwind” one or both factors of $S^{1}$ to factors of $\mathbb{R}$, giving noncompact cosmological spatial slices with topology either $T^{2}\times\mathbb{R}$ or $S^{1}\times\mathbb{R}^{2}$. It is important to note that each of equations (4.1) through (4.6) lift precisely as stated to these “unwound” spacetimes, so some essential features are preserved: namely, the metric is still asymptotic to that of a radiation-dominated FLRW spacetime.
@@
Indeed, the unit timelike vector field $T=(\partial_{\alpha}+r\partial_{r})/(1+r^{2})$ satisfies $T=-\nabla f$. This spacetime is apparently as causally regular as one could hope, asymptotes to a radiation-dominated FLRW geometry, and is geodesically complete. While, as mentioned above, the given metric (4.1) satisfies none of the standard energy conditions in the $r<1$ region, a notable observation is that this is not forced by either Theorem 4.1 or Theorem 4.2, as the noncompact effective cosmological slices at fixed $r>>1$ are not the Cauchy surfaces of this globally hyperbolic spacetime (their past Cauchy horizons are given by the $r=1$ surface of the tubular core). When considering variants of the singularity theorems formulated for globally hyperbolic models of cosmology, then, one must keep in mind that even if the spacetime is globally hyperbolic, the cosmological slices empirically expected to exhibit geodesic convergence may not be among its Cauchy surfaces. In the present example, the foliating paraboloids are “almost” the cosmological slices for very large $r>>1$, but bend away from them near $r\approx 1$.
From the perspective of an initial value problem, an alternative phrasing of these observations is that the $r>>1$ cosmological slices illustrate the existence of initial data sets that are arbitrarily small perturbations of an FLRW initial data set taken in the radiation-dominated regime and that admit a past Cauchy development that is indeed timelike incomplete, as expected from Theorem 4.1, but which is smoothly extendable to a geodesically complete and globally hyperbolic spacetime. Due to the strong curvature nature of the initial singularity in standard FLRW spacetimes, it is often thought that singularities guaranteed by results such as Theorem 4.1 in a perturbation should also be physical: the Eternal Trumpet demonstrates that this may not be the case.
This discussion leads to the natural question of whether one can emulate the Eternal Trumpet’s structure while respecting a physical energy condition:
Question 4.4.
@@
An interesting observation is that, in principle, it seems there is no known obstruction to answering this affirmatively using only radiation as one’s matter source. We remark that in attempting to construct a metric resolving any of Questions 4.3, 4.4, or 4.5, one must keep in mind the constraints imposed by unwinding and rewinding: a metric similar to (4.1) and affirmatively resolving Question 4.5, for example, cannot be so symmetric that it can quotient to $\mathbb{R}^{2}\times T^{2}$, nor can it be so nice that it lifts to a globally hyperbolic metric on $\mathbb{R}^{4}$.
@@ 5 Conclusion
In Section 2, namely Theorem 2.3, we have classified the admissible orientable (not time-orientable) topologies of reflective topological big bangs in $n$ spacetime dimensions as being in one-to-one correspondence with connected, nonorientable (n1)(n-1)-manifolds. For such manifolds which arise as quotients of $\mathbb{R}^{n-1}$ (see Figure 3), these topological big bang models may be endowed with a standard flat FLRW metric (3.1), with only a minor constraint on the scale factor required for smoothness across the reflection. In Section 3, we computed that minimal reflection models do not appear to be able to resolve the horizon problem on their own– a mechanism such as inflation may still be required.
In Section 4.1, we considered a qualitatively different, nonreflective mollification of the singularity in the Itty-Bitty Blender spacetime. This was specified via an explicit metric (4.1) placed on $\mathbb{R}^{2}\times T^{2}$ which is asymptotic to that of a radiation-dominated FLRW geometry, so that the blender’s nontrivial topology may be smoothly adjoined to the radiation-dominated era in standard cosmology. While time-orientable, this spacetime admits closed timelike curves localized in the interior of the blender at the “beginning” of the universe. The nature of the blender automatically resolves the horizon problem, but Hawking’s Theorem 4.2 suggests that the metric cannot be modified to respect the strong energy condition while remaining geodesically complete. An interesting open question, Question 4.3, is whether a modification could respect the more physical dominant energy condition while remaining complete.
In Section 4.2, we discussed that one may avoid both the constraints of certain variants of Hawking’s theorems and any causal pathologies by passing to the universal cover of the blender, the Eternal Trumpet spacetime. This spacetime is endowed with the same metric (4.1), so it is still completely nonsingulary and asymptotic to a radiation-dominated FLRW geometry, but it is also topologically trivial and globally hyperbolic, the strongest available notion of causally well-behaved. The effective cosmological slices in this model provide examples of general relativistic initial data sets which are arbitrarily small perturbations of FLRW initial data sets and which admit a past Cauchy development that is smoothly extendible to a complete and globally hyperbolic spacetime. Known theorems provide no obstruction to there being a modification to the Eternal Trumpet’s metric which respects the dominant energy condition while remaining timelike complete, suggesting Question 4.4. If one seeks to satisfy the strong energy condition as well, known results force one to look between the blender and the trumpet, leading to Question 4.5.
@@ Acknowledgements
The authors would like to thank David Garfinkle for providing helpful and illuminating commentary on the content of this manuscript.

### v3

@@ 3 Beyond Reflection
Throughout Section 2, we have taken some care to emphasize the qualifier reflective in our discussion of topological big bangs. This is because one might imagine that it is possible to replace the singularity of cosmological models with some topology in a manner qualitatively different than Definition 2.1, but still consistent with the broader idea of a topological big bang. We exhibit that this is indeed the case with an explicit example.
@@ 3.1 The Itty-Bitty Blender
Note that the coordinate vector field $\partial_{\beta}$ is always spacelike, so the last $S^{1}$ factor does not interestingly affect the overall structure; we suppress this factor in visualizations and discussion. In contrast, $\partial_{r}$ and $\partial_{\alpha}$ exchange causal character at $r=1$. For $r<1$, $\partial_{r}$ is spacelike and $\partial_{\alpha}$ is timelike: in this range, one should envision that $(r,\theta,\alpha)$ parameterizes a solid toroidal region with $r$ the radial coordinate in the circular cross section at azimuthal angle $\alpha$, and with $\theta$ the angular coordinate around this cross section (see Figure 4).
@@
As in the case of reflective topological big bangs, the causal pathologies of this spacetime (closed timelike curves) are localized to the very early universe. One may still make sense of a classical initial value problem for the future of any spatial $T^{3}$ slice outside the blender (at any $r_{0}>1$), and crucially, an observer in the present has no way of accessing a blender cosmology’s closed timelike curves. It is interesting to note that there is no need for inflation to resolve the horizon problem in this picture: the eternal mixing accommodated by the blender ensures that the whole of spacetime is thoroughly causally embroiled, as every inextendible causal curve emanated from the blender. This resolution is fundamentally distinct from that of inflation, painting a rather different picture of how the early universe may have thermally equilibrated. As pathological as dynamically emerging closed timelike curves would be [16], their existence at the “beginning” of the universe in this manner, perhaps in the regime of quantum gravity and supported by some sort of self-consistent steady-state “swirling” matter distribution, may be permissible.
As discussed in the introduction, we do not attempt here to replicate other successes of inflationary perturbative cosmology (see [1]) or claim that the present model is a genuinely compelling alternative on its own without much more work. Indeed, the metric (3.1) is somewhat contrived and was composed with no thought toward physical energy conditions– in fact, a computation shows that it satisfies none of the standard energy conditions in the $r<1$ region. It is an interesting question whether a qualitatively similar picture can be attained under, say, the dominant energy condition or via a physically motivated matter field; this is discussed further below. Instead, this example, as the whole of this work, is a proof of concept that there exists a rich vein of possibilities for the nature of the early universe beyond the standard discourse.
@@ 3.2 The Blender Unwound: An Eternal Trumpet
Reflective topological big bangs and the Itty-Bitty Blender have an additional out through their causal pathologies– they are clearly not globally hyperbolic, so the cosmological spatial slices $\Sigma_{t}$ are not Cauchy surfaces. While it is clear from their statements, and therefore widely recognized [36], that sufficiently egregious causal pathologies allow one to work around most singularity theorems, the present examples are interesting in that they illustrate that such pathologies can be entirely localized to the universe’s “earliest” moments. It may appear conceivable, then, that one could write down a metric respecting the strong energy condition while having the same qualitative structure as the Itty-Bitty Blender. This is not the case for the blender as we have presented it, with cosmological spatial slices being copies of $T^{3}$ (and hence compact), due to the following variant of Hawking’s theorem (Theorem 55B in [31]; a relevant result is also established in [27]):
@@
Any blender-type construction on $\mathbb{R}^{2}\times T^{2}$, then, cannot respect the strong energy condition without introducing timelike incompleteness. One could reasonably accept this and seek a metric only satisfying an independent (and more physical) energy condition, such as the dominant energy condition [^8] Recall that the dominant energy condition imposes that any four-momentum density seen by observers is future-pointing and causal, a directly physical requirement. In contrast, while the strong energy condition is geometrically convenient, it can be violated by entirely reasonable and simple classical matter models, such as free scalar fields.– there appears to be no obstruction to this in the literature:
Question 3.3.
@@
It would be very interesting for considerations of what constitutes a viable model of cosmology if the answer to this question were affirmative. Another approach to working around the constraints of Theorems 3.1 and 3.2 is to “unwind” one or both factors of $S^{1}$ to factors of $\mathbb{R}$, giving noncompact cosmological spatial slices with topology either $T^{2}\times\mathbb{R}$ or $S^{1}\times\mathbb{R}^{2}$. It is important to note that each of equations (3.1) through (3.6) lift precisely as stated to these “unwound” spacetimes, so some essential features are preserved: namely, the metric is still asymptotic to that of a radiation-dominated FLRW spacetime.
@@
Indeed, the unit timelike vector field $T=(\partial_{\alpha}+r\partial_{r})/(1+r^{2})$ satisfies $T=-\nabla f$. This spacetime is apparently as causally regular as one could hope, asymptotes to a radiation-dominated FLRW geometry, and is geodesically complete. While, as mentioned above, the given metric (3.1) satisfies none of the standard energy conditions in the $r<1$ region, a notable observation is that this is not forced by either Theorem 3.1 or Theorem 3.2, as the noncompact effective cosmological slices at fixed $r>>1$ are not the Cauchy surfaces of this globally hyperbolic spacetime (their past Cauchy horizons are given by the $r=1$ surface of the tubular core). This illustrates that when considering variants of the singularity theorems formulated for globally hyperbolic models of cosmology, then, one must keep in mind that even if the spacetime is globally hyperbolic, the cosmological slices empirically expected to exhibit geodesic convergence may not be among its Cauchy surfaces. In the present example, the foliating paraboloids are “almost” the cosmological slices for very large $r>>1$, but bend away from them near $r\approx 1$.
From the perspective of an initial value problem, an alternative phrasing of these observations is that the $r>>1$ cosmological slices illustrate the existence of initial data sets that are arbitrarily small perturbations of an FLRW initial data set taken in the radiation-dominated regime and that admit a past Cauchy development that is indeed timelike incomplete, as expected from Theorem 3.1, but which is smoothly extendable to a geodesically complete and globally hyperbolic spacetime. That is, one can apply Theorem 3.1 to one of the effective cosmological slices in the Eternal Trumpet spacetime, but the conclusion of incompleteness only applies to its Cauchy development (more precisely, to that portion of the development respecting the strong energy condition), and that development is smoothly extendable to a spacetime which is complete. Due to the strong curvature nature of the initial singularity in standard FLRW spacetimes, it is often thought that “singularities” guaranteed by results such as Theorem 3.1 in a perturbation should also be physical: The Eternal Trumpet demonstrates that this may not be the case.
This discussion leads to the natural question of whether one can emulate the Eternal Trumpet’s structure while respecting a physical energy condition, which would make the observations of the previous paragraph much more compelling:
Question 3.4.
@@
An interesting observation is that, in principle, it seems there is no known obstruction to answering this affirmatively using only radiation as one’s matter source. We remark that in attempting to construct a metric resolving any of Questions 3.3, 3.4, or 3.5, one must keep in mind the constraints imposed by unwinding and rewinding: a metric similar to (3.1) and affirmatively resolving Question 3.5, for example, cannot be so symmetric that it can quotient to $\mathbb{R}^{2}\times T^{2}$, nor can it be so nice that it lifts to a globally hyperbolic metric on $\mathbb{R}^{4}$.
@@ 4 Conclusion
In Theorem 2.3, we have classified the admissible orientable (not time-orientable) topologies of reflective topological big bangs in $n$ spacetime dimensions as being in one-to-one correspondence with connected, nonorientable (n1)(n-1)-manifolds. For such manifolds which arise as quotients of $\mathbb{R}^{n-1}$ (see Figure 3), these topological big bang models may be endowed with a standard flat FLRW metric (2.2), with only a minor constraint on the scale factor required for smoothness across the reflection. In Section 2.2, we computed that minimal reflection models do not appear to be able to resolve the horizon problem on their own– a mechanism such as inflation may still be required.
In Section 3.1, we considered a qualitatively different, nonreflective mollification of the singularity in the Itty-Bitty Blender spacetime. This was specified via an explicit metric (3.1) placed on $\mathbb{R}^{2}\times T^{2}$ which is asymptotic to that of a radiation-dominated FLRW geometry, so that the blender’s nontrivial topology may be smoothly adjoined to the radiation-dominated era in standard cosmology. While time-orientable, this spacetime admits closed timelike curves localized in the interior of the blender at the “beginning” of the universe. The nature of the blender automatically resolves the horizon problem, but Hawking’s Theorem 3.2 suggests that the metric cannot be modified to respect the strong energy condition while remaining geodesically complete. An interesting open question, Question 3.3, is whether a modification could respect the more physical dominant energy condition while remaining complete.
In Section 3.2, we discussed that one may avoid both the constraints of certain variants of Hawking’s theorems and any causal pathologies by passing to the universal cover of the blender, the Eternal Trumpet spacetime. This spacetime is endowed with the same metric (3.1), so it is still completely nonsingular and asymptotic to a radiation-dominated FLRW geometry, but it is also topologically trivial and globally hyperbolic, the strongest available notion of causally well-behaved. The effective cosmological slices in this model provide examples of general relativistic initial data sets which are arbitrarily small perturbations of FLRW initial data sets and which admit a past Cauchy development that is smoothly extendible to a complete and globally hyperbolic spacetime. Known theorems provide no obstruction to there being a modification to the Eternal Trumpet’s metric which respects the dominant energy condition while remaining timelike complete, suggesting Question 3.4. If one seeks to satisfy the strong energy condition as well, known results force one to look between the blender and the trumpet, leading to Question 3.5.
@@ Acknowledgements
The authors would like to thank David Garfinkle for providing helpful and illuminating commentary on the content of this manuscript, as well as the referees for making a number of suggestions improving its clarity and presentation.